Between 2017 and 2018, a small businesspersons income increased from $200,000 to $220,000. The annual state income taxes that she paid increased from $4,30 to $5,100 What was her marginal state income tax rate in 2018? Did this individual experience proportional, progressive, or regressive taxation? Explain briefly This small businesspersons marginal state income tax rate in 2018 was percent.(Enter your response rounded to one decimal place,)
Answer #1

In 2017 : Income = $200,000 and Income tax = $4,300 ,this means tax rate = (4300/200000)100 = 2.15%

In 2018 : Income = $220,000 and Income tax = $5,100, this means tax rate = (5100/220,000)100 = 2.32%

Because as income increases from 2017 to 2018 , income tax rate increases , this implies that the individual experience progressive taxation.

Marginal state income tax rate in 2018 = [(5100-4300)/(220000-200,000)] 100 = (800/20000)100 = 4%